OpenStudy (anonymous):

How do I establish the identity? cos^4x = 1/8(3+4cos(2x) + cos (4x))

OpenStudy (solomonzelman):

Use \(\large\color{black}{ \displaystyle \cos(2x)=2\cos^2x-1 }\)

OpenStudy (anonymous):

ok, how would it work with having 1/8 and 4x

OpenStudy (solomonzelman):

lets just change the angles of (2x) and (4x) into angles of x. (if that makes sense to you)

OpenStudy (solomonzelman):

what will 4cos(2x) be according to this identity ?

OpenStudy (anonymous):

4(2cos^2x-1) ?

OpenStudy (solomonzelman):

yes, and that expands/simplifies to ?

OpenStudy (anonymous):

8cos^2 x-4 ?

OpenStudy (solomonzelman):

yes, there you go

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \cos^4x = \frac{1}{8}(3+4\cos(2x) + \cos (4x)) }\) \(\large\color{black}{ \displaystyle \cos^4x = \frac{1}{8}(3+4(2\cos^2x-1) + \cos (4x)) }\) \(\large\color{black}{ \displaystyle \cos^4x = \frac{1}{8}(3+8\cos^2x-4 + \cos (4x)) }\) \(\large\color{black}{ \displaystyle \cos^4x = \frac{1}{8}(8\cos^2x-1 + \cos (4x)) }\) so far see what is happening?

OpenStudy (anonymous):

sorry i have to leave, hopefully i can finish this up later